OpenStudy (madison5s):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = -16t2 + 576t. After how many seconds does the projectile take to reach its maximum height?

OpenStudy (madison5s):

@Will.H

OpenStudy (will.h):

So we will find the axis of symmetry knowing the it's the x value (time) of the y value (maximum height) Axis of symmetry = -b/2a Axis of symmetry = -576/2(-16) Axis of symmetry = -576/-32 Axis of symmetry = 18

OpenStudy (will.h):

Therefore the time to reach the highest point is 18 seconds

OpenStudy (madison5s):

THANK YOU!!!@Will.H

OpenStudy (will.h):

My pleasure :)